OpenStudy (anonymous):

On the final day of a sale, the price of a rucksack is £12; it's normal price is £32. Write the ratio of the final price to the normal price. Give your answer in it's simplest form.

OpenStudy (watchmath):

32/12=8/3.

OpenStudy (mathmagician):

12/32=3/8

OpenStudy (watchmath):

you were right again mathmagician :). Julia5x takes mathmagician answer!

OpenStudy (anonymous):

thank you both! :)